The temperature of room heated by a heater is 20 degree celcius when outside temperature is -20 degree and it is 10 degree when the outside temperature is -40 degree.The temperature of heater is

Grade:11

1 Answers

Saurabh Koranglekar
3 years ago
the temperature diffrence should be equal for equilibrium
(t−20)=20−(−20);
so,t−10=10−(−40)
t=60